# Meritocracy and Its Discontents: Long-run Effects of Repeated School Admission Reforms

### v1

@@ A.1 The Evolution of the Admission System
Table A.1 shows changes in the number and the admission system of National Higher Schools 1-8 (Kanritsu Koutou Gakkou in Japanese) from 1900 to 1930.313131National Higher Schools were initially called the University Preparatory Division of National Higher School. See Moriguchi (2021) for historical details and the sources of information in Table A.1. Despite the population increase and the growing demand for higher education, the number of Schools 1-8 increased only slightly from 6 to 8 in 1900–1918 due to tight fiscal constraints. With the economic boom of WWI, the government expanded the higher education system and increased the number of Schools from 8 in 1918 to 25 in 1925.323232In the late 1920s, in addition to 25 national higher schools, 3 local public higher schools, 4 private higher schools, and one colonial higher school were established. Such expansions notwithstanding, Schools 1–8 remained the most distinguished among all higher schools throughout the pre-WWII period.
@@ A.3 Data for the Short-run Analysis
To analyze short-run effects of the centralized admission system (Capp), we collect data on application and enrollment outcomes for Schools 1–8 by newly digitizing several administrative and non-administrative sources.
@@ A.3.1 Data on Application Behavior
First, we collect data on the number of applicants by school of their first choice (hereafter first-choice school) for 1900–1930 from the following sources: Ministry of Education Yearbook (Monbushou Nenpou in Japanese) for 1900, 1901, 1908–1916, 1919–1925, and 1928–1930; [^33] Digital images are available online at the National Diet Library Digital Collections. Correspondences from the Ministry of Education to the Tokyo Imperial University for 1903 and 1904; [^34] Digital images are available online at the University of Tokyo Digital Archives. the entrance exam preparation magazine called Middle School World (Chugaku Sekai in Japanese), vol.10, no.12, for 1907; and Higher School Entrance Examination Investigation Report (Koutou Gakkou Nyugaku Shiken ni kansuru Shochousa in Japanese) by the Ministry of Education for 1917, 1918 and 1927; [^35] Digital images are available online at the National Diet Library Digital Collections. no data is available for 1902, 1905, 1906, and 1926. In addition, we collect data on the number of entrants by school for 1900–1930 from Ministry of Education Yearbook.
@@
First-choice school is defined as the school to which an applicant apply under the decentralized admission system and the school which an applicant rank as his first choice under the centralized admission system. There is one issue in this data that requires data cleaning. As Table A.1 indicates, as exceptions under the decentralized system, Schools 7 and 8 are allowed to conduct their entrance exams on different dates from other schools in 1901 (School 7), 1908 (Schools 7 and 8), and 1909–1910 (School 7). In these years, applicants can apply for School 7 or 8 in addition to one of the other schools. For this reason, the number of first-choice applicants to Schools 7 and 8 in these years are double-counted. To correct for double-counting, we estimate the number of first-choice applicants for Schools 7 and 8 in these years using 1) the number of entrants to Schools 7 and 8 in these years and 2) the average ratio of applicants to entrants in 1911–1916 for School 7 and 1909–1916 for School 8, respectively. Using this data, we compute the share of applicants who choose School 1 as their first choice (in Figure 1(a)) and the ratio of first-choice applicants to entrants for each school (in Appendix Figure A.4).
@@
Taking advantage of more detailed data in a supplementary volume of Higher School Entrance Examination Investigation Report of 1917, we collect data on the number of applicants by their first-choice school and by their middle-school prefecture for 1916 and 1917. Middle-school prefecture is defined as the prefecture in which an applicant’s middle school is located. To measure the geographical mobility of applicants, we define application distance (i.e., the distance between an applicant’s middle school and his first-choice school) by the direct (straight-line) distance between the capital of the prefecture in which his middle school is located and the capital of the prefecture in which his first-choice school is located. The distance data is provided by the Geospatial Information Authority of Japan. These data is used to examine the impact of centralized admissions on (a) the share of applicants whose first choice is School 1 and (b) the application distance by regions (Appendix Table A.3).
@@ A.3.2 Data on Enrollment Outcomes
To analyze enrollment outcomes, we use Higher School Student Registers (Gakkou Ichiran in Japanese) published annually by Schools 1–8.363636Digital images are available online at the National Diet Library Digital Collections. For each school, as a proxy for the number of entrants, we collect data on the number of first-year students in the university preparatory course (daigaku yoka in Japanese). Our data starts in 1896 for Schools 1, 2, 4, and 5, 1897 for School 3, 1900 for School 6, 1901 for School 7, and 1908 for School 8, reflecting the year of establishment for each school (School 3 was established in 1896, but the university preparatory course started in 1897). [^37] The data is missing in the following years: 1929 for School 3, 1904 and 1907 for School 5, 1920 for School 6, and 1909 for School 7. Birth prefecture is the prefecture of a student’s legal domicile (honsekichi in Japanese) recorded in the official family registry system (koseki in Japanese). We include students born in any prefectures, but exclude foreign-born students and students born in colonies.
@@
To measure the geographical mobility of entrants in Figure 1(b), we define enrollment distance (i.e., the distance between an entrant’s birth prefecture and the school he entered) by the direct distance between the capital of the birth prefecture and the capital of the prefecture in which the school he entered is located. The distance data is provided by the Geospatial Information Authority of Japan. To produce Figure 1(b), we focus on the sample of schools that use single-school application under the decentralized system. As Schools 7 and 8 are allowed to conduct their entrance exams on different dates from other schools in 1901 (School 7), 1908 (Schools 7 and 8), and 1909–1910 (School 7), applicants can apply for these schools in addition to one of the rest of the schools in these years even under the decentralized system. We exclude these schools in these years from the sample, since the feature of multiple-school application contaminates the essential aspect of the decentralized system.
@@
To analyze the impact of centralized admissions on the geographical composition of entrants (Figure 2(a), Figure 2(b), and Table 1), we control for the size of potential applicants and the number of competing schools in each prefecture, using the following variables. From Ministry of Education Yearbook, we collect data on the number of middle-school graduates (including both public and private schools) by prefecture (defined by school location)  [^38] The data on middle-school graduates is missing for 1920 and is linearly interpolated. as well as the number of higher schools (including national, public, and private schools) by prefecture (defined by school location) for 1900–1930. The total number of higher schools increased from 8 (consisting of 8 national schools) in 1918 to 32 (consisting of 25 national, 3 public, and 4 private schools) in 1930. The Tokyo area is defined by seven prefectures located within 100 km from Tokyo (i.e., Chiba, Gumma, Ibaraki, Kanagawa, Saitama, Tochigi, and Tokyo prefectures) measured by the distance between two prefectural capitals (Appendix Table A.3).
@@
Finally, to test the exogeneity of the timing of admission reforms, in addition to the data collected above, we collect data on the mean age of entrants for 1905–1930 from Ministry of Education Yearbook. [^39] No data on the mean age of entrants is available for 1900–1904. We also collect data on government expenditures on national higher education (the sum of ordinary and extraordinary expenditures spent on National Higher Schools and Imperial Universities) for 1900–1930 from Ministry of Education Yearbook. Our analysis is presented in Appendix Table A.7.
@@ A.4 Data for the Long-run Analysis
To assess long-run effects of the centralized admission system (Capp), we provide two sets of empirical analyses. In the first analysis, we compile prefecture-cohort level data of career elites and compare career outcomes of urban- and rural-born individuals across birth cohorts that differed in their exposure to Capp. In the second analysis, we construct cohort level data of central government officials using the complete list of individuals who have passed the Higher Civil Service Examinations and compare promotion outcomes of these individuals across cohorts that differed in their exposure to Capp.
A.4.1 Data on the Regional Distribution of Career Elites
@@
To measure students’ career outcomes, the first long-run analysis uses Personnel Inquiry Records (Jinji Koushin-roku in Japanese), which compiles a highly selective list of “socially distinguished individuals” (shakai-teki meishi in Japanese) and provides their biographical information. We use the 1939 edition of the PIR in our main analysis and the 1934 edition for a supplementary analysis. [^40] Digital images of the PIR are available online at the National Diet Library Digital Collections. We thank Hidehiko Ichimura and Yasuyuki Sawada for sharing their digitized PIR data.
@@
In selecting the socially distinguished individuals, the PIR refers to multiple sources such as the government personnel directory (Shokuin-roku in Japanese), the directories of banks and companies (Teikoku Ginkou Kaisha Youroku and Zenkoku Ginkou Kaisha Youroku in Japanese), the directory of the national chamber of commerce and industry members (Zenkoku Shoukou Kaigisho Giin Meibo in Japanese), and the directory of Japanese notables (Nihon Shinshi-roku in Japanese) (see PIR 1934, p.2, for details). [^41] Digital images of these directories are available online at the National Diet Library Digital Collections.
@@
The government personnel directory provides a complete list of public servants (with their job titles and ranks) in national and local governments, including the national diet members, civil and military officials, and Imperial University professors. One of the directories of banks and companies provides a complete list of directors of banks and companies whose capital is 300,000 JPY or above (Teikoku Ginkou Kaisha Youroku 1938, p.2). The directory of Japanese notables includes “wealthy persons” (shisanka in Japanese, defined by individuals who paid income tax of 80 JPY or greater or business tax of 70 JPY or greater) living in urban areas in 24 prefectures (out of 47 prefectures) in Japan (Nihon Shinshi-roku, 1938, p.i).
@@
We use the following information from the PIR for each individual: full name, birth year, birth prefecture, the prefecture of residence, final education, occupation titles and organization names, corporation type (if the organization is incorporated), the medal for merit and the court rank awarded (if any), and the amounts of national income tax and business tax paid in the previous year (if any). To check for accuracy, the PIR verifies the information of birth date and birth prefecture for each individual by obtaining a transcript of the official family register (PIR 1934, p.2).
Defining the Groups of Career Elites
@@
As described in the main text, we define (mutually non-exclusive) groups of elites among the individuals listed in the PIR. We also define landlords as individuals whose occupations include landlord, but excluding the top 0.05% income earners, prestigious medal recipients, corporate executives, top politicians and bureaucrats, and Imperial University professors. To determine individual’s occupations, we check if one’s occupational titles and organization names contain specific Japanese characters that signify a given occupation. The precise definitions of corporate executives, top politicians and bureaucrats, Imperial University professors, landlords, scholars, engineers, physicians, and lawyers are provided in Appendix Table A.8.
@@
Third, to control for local demographic changes, we estimate the prefecture-level male birth population of the cohorts born in 1874–1894 using the following data sources: (a) Japanese Population Census (Nihon Zenkoku Kokou-hyo in Japanese) in 1880–1892 that provides the number of male births by prefecture in each year and (b) Japanese Imperial Population Census (Nihon Teikoku Minseki Kokou-hyou in Japanese) in 1886 that provides the age-specific population by prefecture (from which we use the population of males who were born in 1874–1885 and were 1–12 years old in 1886). [^43] Digital images of population censuses are available online at the National Diet Library Digital Collections. For the cohorts born in 1874–1879, we estimate their birth population combining (a) and (b) as follows. For the cohorts born in 1880–1886, we define the survival rate of cohort $c$ up to 1886 in prefecture $j$ by the number of age-specific population of cohort $c$ in 1886 from (b) divided by the number of births of cohort $c$ from (a). We then estimate the following equation using OLS:
@@
Estimating the Intensity of Exposure to Capp
@@
Before 1918, the qualifications for the higher school entrance exam were male middle-school graduates who were at least 17 years old. Accordingly, applicants typically took the exam for the first time at age 17. Given this, in the regression analyses in Table 2 Panels A–C, Table A.14, and Table 3, we simply assume that, among cohorts who were born in 1880–1894 and turned age 17 in 1897–1911, only the cohorts who turned age 17 during the first period of centralized admissions (Capp) in 1902–1907 were fully exposed to Capp, while the rest of the cohorts were fully exposed to Dapp.
@@
Precisely speaking, however, a nontrivial number of unsuccessful applicants retook the exam multiple times in the subsequent years at age greater than 17. For this reason, we also estimate each cohort’s intensity of exposure to Capp that takes into account the presence of exam re-takers as follows. We use a special table in Government Gazette No.5838 (published in December 17, 1902) that reports the number of applicants and the number of admitted applicants in 1902 by the year of middle-school graduation. According to the table, 63.9% of the applicants in 1902 graduated from middle schools in 1902 (the same year), 27.3% in 1901 (the previous year), 7.2% in 1900 (two years before), 1.3% in 1899 (three years before), 0.2% in 1898 (four years before), and 0.1% in 1897 (five years before).
@@
We use the 1902 data to compute the total number of exams taken and the number of exams taken under Capp for each cohort, where cohort is defined by the year of middle-school graduation. This exercise requires us to estimate two sets of parameters. The first set of parameters is the probability of passing the exam in year $t$ for applicants in cohort $c$ who took the exam $n$ times before year $t$, denoted by $a_{n}^{t}$ (where $n=0,...,5$ and $t=1896,...,1911$). Note that the subscript for cohort $c$ is suppressed because $c=t-n$. The second set of parameters is the probability of applicants (who took the exam $n$ times) retaking the exam in year $t+1$ conditional on failing the exam in year $t$, denoted by $r_{n}^{t}$. Because we only have the data for a single year, we assume that these parameters are constant across years (and therefore expressed as $a_{n}$ and $r_{n}$) and that the cohort size is constant across cohorts. Under these assumptions, the probability of passing the exam is estimated to be $a_{0}=0.352$, $a_{1}=0.377$, $a_{2}=0.313$, $a_{3}=0.286$, $a_{4}=0.429$, $a_{5}=0.200$. The probability of retaking the exam is estimated to be $r_{0}=0.660$, $r_{1}=0.423$, $r_{2}=0.255$, $r_{3}=0.175$, $r_{4}=0.011$, $r_{5}=0.000$.
@@
Using these parameters, we simulate the total number of exams taken by each cohort and compute the share of the exams taken under Capp, which we define as the intensity of exposure to Capp. Here we assume that all cohorts graduated from middle-schools at age 17 (so that birth cohorts and cohorts defined by the year of middle-school graduation coincide). The estimated intensity of exposure to Capp is 0.00, 0.02, 0.09, 0.36, 1.00, 0.98, 0.91, 0.64, and 0.00 for the cohort who turned age 17 in 1896–98, 1899, 1900, 1901, 1902–04, 1905, 1906, 1907, and 1908–11, respectively. We color cohorts according to their intensity of exposure to Capp in Figure 3 and Appendix Figures A.7, A.10, and A.11. We also use the estimated intensity of exposure in the regression analyses in Table 2 Panel D and Appendix Table A.17.
A.4.2 Data on the National Production of Career Elites

### v2

@@ A.1 The Evolution of the Admission System
Table A.1 shows changes in the number and the admission system of Schools 1–8 (Kanritsu Koutou Gakkou in Japanese) from 1900 to 1930. Despite the population increase and the growing demand for higher education, the number of National Higher Schools increased only slightly from 6 to 8 in 1900–1918 due to tight fiscal constraints. With the economic boom of WWI, the government expanded the higher education system and increased the number of Schools from 8 in 1918 to 25 in 1925.333333In the late 1920s, in addition to 25 national higher schools, 3 local public higher schools, 4 private higher schools, and one colonial higher school were established. Such expansions notwithstanding, Schools 1–8 remained the most distinguished among all higher schools throughout the pre-WWII period.
@@ A.3 Data for the Short-run Analysis
To analyze short-run effects of the centralized admission system, we collect data on application and enrollment outcomes for Schools 1–8 by newly digitizing several administrative and non-administrative sources.
@@ A.3.1 Data on Application Behavior
We collect data on the number of applicants by school of their first choice (hereafter first-choice school) for 1900–1930 from the following sources: the Ministry of Education Yearbook (Monbushou Nenpou in Japanese) for 1900, 1901, 1908–1916, 1919–1925, and 1928–1930; [^34] Digital images are available online at the National Diet Library Digital Collections. Correspondences from the Ministry of Education to the Tokyo Imperial University for 1903 and 1904; [^35] Digital images are available online at the University of Tokyo Digital Archives. the entrance exam preparation magazine called Middle School World (Chugaku Sekai in Japanese), vol.10, no.12, for 1907; and Higher School Entrance Examination Investigation Report (Koutou Gakkou Nyugaku Shiken ni kansuru Shochousa in Japanese) by the Ministry of Education for 1917, 1918 and 1927; [^36] Digital images are available online at the National Diet Library Digital Collections. no data is available for 1902, 1905, 1906, and 1926. In addition, we collect data on the number of entrants by school for 1900–1930 from the Ministry of Education Yearbook.
@@
First-choice school is defined as the school to which an applicant applies under the decentralized admission system and the school which an applicant ranks as his first choice under the centralized admission system. There is one issue in this data that requires data cleaning. As Table A.1 indicates, as exceptions under the decentralized system, Schools 7 and 8 are allowed to conduct their entrance exams on different dates from other schools in 1901 (School 7), 1908 (Schools 7 and 8), and 1909–1910 (School 7). In these years, applicants can apply for School 7 or 8 in addition to one of the other schools. For this reason, the number of first-choice applicants to Schools 7 and 8 in these years are double-counted. To correct for double-counting, we estimate the number of first-choice applicants for Schools 7 and 8 in these years using 1) the number of entrants to Schools 7 and 8 in these years and 2) the average ratio of applicants to entrants in 1911–1916 for School 7 and 1909–1916 for School 8, respectively. Using this data, we compute the share of applicants who choose School 1 as their first choice (in Figure 1(a)) and the ratio of first-choice applicants to entrants for each school (in Appendix Figure A.4).
@@
Taking advantage of more detailed data in a supplementary volume of the Higher School Entrance Examination Investigation Report of 1917, we collect data on the number of applicants by their first-choice school and by their middle-school prefecture for 1916 and 1917. Middle-school prefecture is defined as the prefecture in which an applicant’s middle school is located. To measure the geographical mobility of applicants, we define application distance (i.e., the distance between an applicant’s middle school and his first-choice school) by the direct (straight-line) distance between the capital of the prefecture in which his middle school is located and the capital of the prefecture in which his first-choice school is located. The distance data is provided by the Geospatial Information Authority of Japan. These data is used to examine the impact of centralized admissions on (a) the share of applicants whose first choice is School 1 and (b) the application distance by regions (Appendix Table A.3).
@@ A.3.2 Data on Enrollment Outcomes
To analyze enrollment outcomes, we use Higher School Student Registers (Gakkou Ichiran in Japanese) published annually by Schools 1–8.373737Digital images are available online at the National Diet Library Digital Collections. For each school, as a proxy for the number of entrants, we collect data on the number of first-year students in the university preparatory course (daigaku yoka in Japanese). Our data starts in 1896 for Schools 1, 2, 4, and 5, 1897 for School 3, 1900 for School 6, 1901 for School 7, and 1908 for School 8, reflecting the year of establishment for each school (School 3 was established in 1896, but the university preparatory course started in 1897). [^38] The data is missing in the following years: 1929 for School 3, 1904 and 1907 for School 5, 1920 for School 6, and 1909 for School 7. Birth prefecture is the prefecture of a student’s legal domicile (honsekichi in Japanese) recorded in the official family registry system (koseki in Japanese). We include students born in any prefecture, but exclude foreign-born students and students born in colonies.
@@
To measure the geographical mobility of entrants in Figure 1(b), we define enrollment distance (i.e., the distance between an entrant’s birth prefecture and the school he entered) by the direct distance between the capital of the birth prefecture and the capital of the prefecture in which the school he entered is located. The distance data is provided by the Geospatial Information Authority of Japan. To produce Figure 1(b), we focus on the sample of schools that use single-school application under the decentralized system. As Schools 7 and 8 are allowed to conduct their entrance exams on different dates from other schools in 1901 (School 7), 1908 (Schools 7 and 8), and 1909–1910 (School 7), applicants can apply for these schools in addition to one of the rest of the schools in these years even under the decentralized system. We exclude these schools in these years from the sample.
@@
To analyze the impact of centralized admissions on the geographical composition of entrants (Figure 2(a), Figure 2(b), and Table 1), we control for the size of potential applicants and the number of competing schools in each prefecture, using the following variables. From the Ministry of Education Yearbook, we collect data on the number of middle-school graduates (including both public and private schools) by prefecture (defined by school location) [^39] The data on middle-school graduates is missing for 1920 and is linearly interpolated. as well as the number of higher schools (including national, public, and private schools) by prefecture (defined by school location) for 1900–1930. The total number of higher schools increased from 8 (consisting of 8 national schools) in 1918 to 32 (consisting of 25 national, 3 public, and 4 private schools) in 1930.
@@
Finally, to test the exogeneity of the timing of admission reforms, in addition to the data collected above, we collect data on the mean age of entrants for 1905–1930 from the Ministry of Education Yearbook. [^40] No data on the mean age of entrants is available for 1900–1904. We also collect data on government expenditures on national higher education (the sum of ordinary and extraordinary expenditures spent on National Higher Schools and Imperial Universities) for 1900–1930 from the Ministry of Education Yearbook. Our analysis is presented in Appendix Table A.7.
@@ A.4 Data for the Long-run Analysis
A.4.1 Data on the Regional Distribution of Elites
@@
To measure students’ career outcomes, the first long-run analysis uses Personnel Inquiry Records (Jinji Koushin-roku in Japanese), which compiles a selective list of “socially distinguished individuals” (shakai-teki meishi in Japanese) and provides their biographical information. We use the 1939 edition of the PIR in our main analysis and the 1934 edition for a supplementary analysis. [^41] We thank Hidehiko Ichimura and Yasuyuki Sawada for sharing their digitized the 1934 and 1939 PIR data.
@@
In compiling a list of socially distinguished individuals, the PIR explicitly states that it used multiple sources, including the government personnel directory (Shokuin-roku in Japanese), the directories of banks and companies (Teikoku Ginkou Kaisha Youroku and Zenkoku Ginkou Kaisha Youroku in Japanese), the directory of the national chamber of commerce and industry members (Zenkoku Shoukou Kaigisho Giin Meibo in Japanese), and the directory of Japanese notables (Nihon Shinshi-roku in Japanese) (PIR (1934), p.2). [^42] Digital images of these materials are available online at the National Diet Library Digital Collections.
@@
The government personnel directory provides a complete list of public servants (with their job titles and ranks) in national and local governments, including the Imperial Diet members, civil and military officials, and Imperial University professors. The directory of banks and companies provides a complete list of directors of banks and companies whose capital is 300,000 JPY or above (Teikoku Ginkou Kaisha Youroku (1938), p.2). The directory of Japanese notables includes “wealthy persons,” defined by individuals who paid income tax of 80 JPY or greater or business tax of 70 JPY or greater, living in urban areas in 24 prefectures (out of 47 prefectures) in Japan (Nihon Shinshi-roku (1938), p.i).
@@
We use the following information from the PIR for each individual: full name, birth year, birth prefecture, the prefecture of residence, final education, occupation titles and organization names, corporation type (if the organization is incorporated), the medal for merit and the court rank awarded (if any), and the amounts of national income tax and business tax paid in the previous year (if any). To check for accuracy, the PIR verifies the information of birth date and birth prefecture for each individual by obtaining a transcript of the official family register (PIR (1934), p.2).
Defining the Groups of Elites
@@
As described in the main text, we define (mutually non-exclusive) subgroups of elites among the individuals listed in the PIR. To determine an individual’s occupations, we search if one’s occupational titles and organization names contain specific Japanese characters that signify a given occupation. The precise definitions of corporate executives, top politicians and bureaucrats, Imperial University professors, landlords, scholars, engineers, physicians, and lawyers are provided in Appendix Table A.9.
@@
Finally, we define all occupational elites as all individuals listed in the PIR except for those individuals whose only occupation is peerage membership or landlord. We exclude these elites because they became elites typically by inheritance.
@@
Third, to control for local demographic changes, we estimate the prefecture-level male birth population of the cohorts born in 1874–1894 using the following data sources: (a) Japanese Population Census (Nihon Zenkoku Kokou-hyo in Japanese) in 1880–1892, which provides the number of male births by prefecture in each year, and (b) Japanese Imperial Population Census (Nihon Teikoku Minseki Kokou-hyou in Japanese) in 1886, which provides the age-specific population by prefecture (from which we use the population of males who were born in 1874–1885 and were 1–12 years old in 1886). [^44] Digital images of population censuses are available online at the National Diet Library Digital Collections. For the cohorts born in 1874–1879, we estimate their birth population combining (a) and (b) as follows. For the cohorts born in 1880–1886, we define the survival rate of cohort $c$ up to 1886 in prefecture $j$ by the number of age-specific population of cohort $c$ in 1886 from (b) divided by the number of births of cohort $c$ from (a). We then estimate the following equation using OLS:
@@
Estimating the Intensity of Exposure to Centralization
@@
Before 1918, the qualifications for the higher school entrance exam were male middle-school graduates who were at least 17 years old. Accordingly, applicants typically took the exam for the first time at age 17. Given this, in the regression analyses in Table 2 Panels A–C, Table A.15, and Table 3, we simply assume that, among cohorts who were born in 1880–1894 and turned age 17 in 1897–1911, only the cohorts who turned age 17 during the first period of centralized admissions in 1902–1907 were fully exposed to centralization, while the rest of the cohorts were fully exposed to decentralization.
@@
Precisely speaking, however, a nontrivial number of unsuccessful applicants retook the exam multiple times in the subsequent years at age greater than 17. For this reason, we also estimate each cohort’s intensity of exposure to centralization by taking into account the presence of exam re-takers as follows. We use a special table in the Government Gazette No.5838 (published in December 17, 1902) that reports the number of applicants and the number of admitted applicants in 1902 by the year of middle-school graduation. According to the table, 63.9% of the applicants in 1902 graduated from middle schools in 1902 (the same year), 27.3% in 1901 (the previous year), 7.2% in 1900 (two years before), 1.3% in 1899 (three years before), 0.2% in 1898 (four years before), and 0.1% in 1897 (five years before).
@@
We use the 1902 data to compute the total number of exams taken and the number of exams taken under centralized admissions for each cohort, where cohort is defined by the year of middle-school graduation. This exercise requires us to estimate two sets of parameters. The first set of parameters is the probability of passing the exam in year $t$ for applicants in cohort $c$ who took the exam $n$ times before year $t$, denoted by $a_{n}^{t}$ (where $n=0,...,5$ and $t=1896,...,1911$). Note that the subscript for cohort $c$ is suppressed because $c=t-n$. The second set of parameters is the probability of applicants (who took the exam $n$ times) retaking the exam in year $t+1$ conditional on failing the exam in year $t$, denoted by $r_{n}^{t}$. Because we only have the data for a single year, we assume that these parameters are constant across years (and therefore expressed as $a_{n}$ and $r_{n}$) and that the cohort size is constant across cohorts. Under these assumptions, the probability of passing the exam is estimated to be $a_{0}=0.352$, $a_{1}=0.377$, $a_{2}=0.313$, $a_{3}=0.286$, $a_{4}=0.429$, $a_{5}=0.200$. The probability of retaking the exam is estimated to be $r_{0}=0.660$, $r_{1}=0.423$, $r_{2}=0.255$, $r_{3}=0.175$, $r_{4}=0.011$, $r_{5}=0.000$.
@@
Using these parameters, we simulate the total number of exams taken by each cohort and compute the share of the exams taken under centralized admissions, which we define as the intensity of exposure to centralization. Here we assume that all cohorts graduated from middle-schools at age 17 (so that birth cohorts and cohorts defined by the year of middle-school graduation coincide). The estimated intensity of exposure to centralization is 0.00, 0.02, 0.09, 0.36, 1.00, 0.98, 0.91, 0.64, and 0.00 for the cohort who turned age 17 in 1896–98, 1899, 1900, 1901, 1902–04, 1905, 1906, 1907, and 1908–11, respectively. We color cohorts according to their intensity of exposure to centralization in Figure 3 and Appendix Figures A.7, A.10, and A.11. We also use the estimated intensity of exposure in the regression analyses in Table 2 Panel D and Table 5.
A.4.2 Data on the National Production of Top-Ranking Higher Civil Officials